# Removing Reflections from RAW Photos

### v2

@@ D.1 Mixture search
Well mixed $m$ are identified by computing the SSIM between $m$ and $t$ as a block-wise image, and checking if the mean of this SSIM image is within a useful range: if the SSIM is too high, the reflection is imperceptible; if it is too low, the mixture is not visually interpretable, even by a human. We compute this single channel SSIM image as a weighted average of the corresponding per-channel SSIM images. The weights are the average value in each color channel, which better accounts for strongly colored images. Lastly, the standard deviation of this single channel SSIM image is checked to remove reflections that are imperceptible, but nonetheless produce a low mean SSIM by spreading their power broadly (they have low spatial variance).
@@ E.1 Base model
The base model is designed to leverage local and global features, Fig. 2, and produce $256^{2}$ pixel outputs in about $1$ second on a mobile device to meet req. 4 (see Sec. 1).
@@
The contextual image, $c$, is passed through the same F-Net and P-Net using the same weights as $m$. The features of $c$ and $m$ at the lowest resolution are input to an affine prediction module, Fig. 2 (lower right). This module first vectorizes its two $64\times 4\times 4$ inputs, and passes them through a fully connected layer to transform them into two $64$-$\mathrm{D}$ vectors. These vectors embody $64$ pairs of channels, which we concatenate and input to an MLP (labeled paired-channel MLP) that predicts affine transforms that modify the features of the FPN during the final low-to-high pass.
The paired-channel MLP is a series of grouped convolutions that implement $64$ independent MLPs followed by a fully connected layer. These $64$ MLPs each have $2$ inputs, $2$ hidden, and $1$ output dimension, with leaky ReLUs after each layer. The inputs to these MLPs are corresponding pairs of channels from $m$ and $c$. The outputs compose a single $64$-$\mathrm{D}$ vector that is input to a fully connected layer to predict $64\times K\times 2$ affine transforms, two for each of the $64\times K$ channels and levels of the FPN. [^4] Note that we include two levels at $256^{2}$ pixels, in correspondence with the resolutions of the features that we extract from the backbone. Conceptually, this paired-channel MLP has the capacity to compare $c$ and $m$ to identify channels that match with the reflection scene, and to determine how to transform those channels to aid reflection removal.
The predicted affines from the paired-channel MLP are used in the final low-to-high pass of the FPN to perform a series of modulated merge operations, Fig. 2 (upper right). These merge operations use two affine transforms per feature channel, labeled $a$ and $b$ in Fig. 2. Transforms $a$ are used to perform a conv-mod-deconv operation ala StyleGan [23] on the features of $m$ from FPN level $K$. These features are subsequently combined with the features from level $K+1$ by resampling the latter features 2×2$\times$2 × and using fast normalized fusion [45] (labeled FPN Combine). These combined features are modified with a second conv-mod-deconv operation using the second group of affines, $b$.
The final modulated merge produces $64$-$\mathrm{D}$ features at $256^{2}$ pixels. These features are concatenated with $m$ and the features from the first layer, for a total of $3+32+64=99$ channels. A convolutional finishing module is then applied. This module has the capacity to further identify and finally render $t$ and $r$. To simplify comparison to prior work, our finishing module is the head in [58]. The first layer is $1\times 1$, and projects the $99$ input features to $64$ dimensions, which are maintained throughout the remaining operations. Those operations are $3\times 3$ convolutions at dilation rates $(1,2,4,8,16,32,64,1)$, each followed by a batch norm and leaky ReLU. A final $1\times 1$ convolution generates $6$ channels: $t$ and $r$.
@@ E.2 Upsampler
The upsampler architecture is summarized in Sec. 4.2 and Fig. 3. It transforms low resolution outputs $t$, $r$ from the base model to a flexible output resolution. We use a Gaussian pyramid and apply the same upsampler at each level, Fig. 3. The upsampler projects the low resolution, $3$-channel inputs $(m,r,t)$ into a high dimensional space $\phi$ using convolutions in an expand-and-contract pattern. We use $3\times 3$ kernels for feature expansion, and $1\times 1$ kernels for contraction. There are $3$ layers with hidden dimensions $(32,16),(64,32),(128,64)$. We use leaky ReLU nonlinearities between the hidden layers, and no skip connections. Batch norms are omitted to facilitate a feature matching process, described next.
The components $t$ and $r$ are separated by identifying low resolution features $\phi_{t}$ and $\phi_{r}$ in the low resolution mixture $\phi_{m}$. We predict $2$, per-pixel, per-channel low resolution masks using a mask prediction module, Fig. 3 (bottom), which uses a paired-channel MLP (defined in Sec. E.1) to predict its affine transforms (see also Sec. 4.2). The joint mask predictor also uses a paired-channel MLP, but it directly outputs the final masks rather than affine transforms. The input, hidden, and output dimensions of both paired-channel MLPs are $(2,2,2,1,2)$. The final layer is fully connected. As noted in Sec. E.1, these MLPs can be implemented efficiently as a series of grouped $1\times 1$ convolutions.
@@ F.1 Evaluation methods
Ground truth capture. Mixture images $m$ were captured with ground truth $r$ by placing a black material behind the glass and taking a second photo. Images $t$ were computed $t=m-r$ in linear sRGB after normalizing the exposures and using the white point of $m$ (see ACR Step 6, Sec. 3). Nonetheless, we found it necessary to capture $m$ and $r$ with fixed exposures and white points to avoid imprecisions in the values that are stored in RAW metadata.
@@
Ground truth photographs. Ground truth images were captured in three common scenarios: 1) looking out of a home window, Fig. 9; 2) photographing artwork, Fig. A12 (left); and 3) looking into a display case, Fig. A12 (right). In scenario one, the illuminant in the reflection scene is approximately $3$,$000\mathrm{K}$, and the white point of the mixture image is $7$,$300\mathrm{K}$. In scenario two, these values are $6$,$000\mathrm{K}$ and $6$,$850\mathrm{K}$, and in scenario three they are $6$,$000\mathrm{K}$ and $3$,$627\mathrm{K}$.
Figure A12: Comparisons with ground truth at $256\times 384$ and $2048\times 3072$. The re-trained, low-resolution methods Zhang [58] and DSRNet [20] are upsampled using our upsampler, and both low- and high-resolution results are shown. The SSIM of the predicted $t$ and $r$ is reported at low resolution (labeled t, r) and high (labeled T, R). Errors in ground truth are outlined and omitted from the SSIM.
@@ F.2 Base model comparisons
Reflections with ground truth. In Fig. A12 we show two images with ground truth reflections: photographing artwork, and looking into a display case. These complement Fig. 9, in which a photo was taken when looking out of a home window. The home window view includes an interior reflection that is strongly yellow due to the indoor illuminant color. In contrast, the artwork in Fig. A12 is illuminated by the same light source as the reflection scene, which produces a correctly white balanced reflection that consequently has more diverse colors. In contrast, the display case in Fig. A12 (right) reflects an outdoor scene with a very different white balance, which produces a reflection with a strong blue color cast. This outdoor reflection is visible over broad regions because the illuminant is powerful enough to reflect off of the diffuse ground and sidewalk surfaces with sufficient intensity to be clearly visible over the contents of the display case. These exterior reflections are also qualitatively different from those in photo of the artwork, where the reflection is sparse. The SSIM of the artwork is therefore high on average ($0.994$), but the reflections are locally strong, whereas the SSIM of the display case is low ($0.833$) because the reflection affects broad regions.
@@
Our base models improve the SSIM of $t$ and $r$ in all of these ground truth examples (labeled in lower case t, r), and this extends to the upsampled results (labeled in uppercase T, R) whereas prior works do not perform as well (Fig. A12 and Fig. 9). Our contextual model produces a more correct transmission and reflection image on the artwork. On the display case, the contextual model improves the reflection, whereas the cars are not fully removed from the transmission. We believe this variation results from saturated regions in the sky of the contextual photo, where we use the as-shot illuminant color in the EXIF to recover the color of the saturated pixels. Lastly, the method of Zhang [58] associates blue colored content with the reflection in both of these images, but this is incorrect for the painting; the transmission image is consequently distorted.
@@
Figure A13: Results for base models at $256^{2}$ pixels. Reflections marked “+X” are brightened by X stops compared to the transmission.
@@
Figure A14: Results for base models at $256^{2}$ pixels. Reflections marked “+X” are brightened by X stops compared to the transmission.
@@
We found that our model removes blue colored reflections consistently well, and we believe that this results from their commonness (outdoor illuminants are powerful and therefore frequently create reflections that appear blue when mixed with interior scenes). The yellow color of interior reflections on outdoor scenes also seems to help, as our model can separate textured objects like the painting on the wall in Fig. 9 from the tree texture. We also tried illuminating the indoor scene in Fig. 9 with a special studio light that is much more powerful than a typical interior light source. This created a warm indoor reflection that was analogous to the display case in Fig. A12 (right), where the reflection covers large parts of the transmission scene. Our model results are less consistent in this artificial situation. We believe this is because it is rare for interiors to be flooded with such strong lights, and so their reflections are uncommon in our training data. We find that interior reflections tend to appear in small regions because most artificial light sources are weak—only objects near the light will be bright enough to reflect over outdoor scenes. At nighttime, however, consumer illuminants easily reflect over dark cityscapes. We found that our model results are less consistent on such images. This can be improved by augmenting the dataset with source images $t=T(i)$ that were taken at nighttime.
@@
Reflections in the wild. In Fig. A13, A14, and Fig. 9 we show results on reflections that were captured in-the-wild, where it was not possible to capture ground truth: (left) shopping in the morning, (middle) looking into a building at midday in a city, and (right) photographing artwork in an outdoor mall (Fig. A13); while traveling and photographing artwork from a city street (A14). We compare to Zhang [58] and DSRNet [20] using their published $8$-bit models (bottom two rows). The $8$-bit models do not consistently remove in-the-wild reflections, with the exception that the method of Zhang et al. seems to have learned to remove blue colored content (left); see also Fig. A15 (top). We retrained these prior works on our photometrically accurate training data (marked “retrained”), and they improve significantly. This suggests that the muted response of the $8$-bit models on in-the-wild reflections results from differences between their training data and real world reflections, and furthermore that the training process of prior works is insufficient: pre-training on photometrically inaccurate images, and fine-tuning on small datasets of ground truth reflections does not produce models that generalize as well. Looking closely, however, note that the retrained models do still introduce blur and colored artifacts.
@@
Our methods perform well across these diverse in-the-wild use cases. In Fig. A13, both our contextual and non-contextual models separate the strongly blue colored reflection, the complex cityscape reflection, and the artwork reflection (in the entrance to an indoor shopping mall). The contextual model improves each of these cases, excepting the transmission image for the shopping photo. We believe this also results from the saturated regions in the upper right of the contextual photo, where we have used the as-shot illuminant color in the EXIF to recover the pixels.
@@
Subtle reflections. In Fig. A15 we show results on images with subtle reflections, which are common in art museums where the glass and lighting are optimized to reduce reflections. Our models are able to remove these reflections, and recover $r$ even when it is invisible to the naked eye. In the painting (Rembrandt’s “St. Matthew and the Angel”), the recovered reflection has a strong blue color
Figure A15: Reflection recovery. Our model separates reflections that can be difficult to spot with the naked eye.
@@
due to the special glass that is used in museums, and it correctly depicts the photographer and gallery. The colors in the recovered painting are also more correct. The specular reflection from the surface of the painting (top) is not removed, and we believe that this is the desired result. We have found that our models will sometimes remove sharp specular regions that are similar in color to the reflection from the glass, and this is visible in the upper quarter of the reflection image, where some specular texture from the surface of the painting is visible, and has been removed from the transmission.
Figure A16: Removing lens flares. Although our training data do not include images of lens flares, our model can sometimes remove them.
Figure A17: Upsampler performance comparison. We upsample a ground truth reflection image from $256\times 384$ to $2048\times 3072$ using our method and V-DESIRR [36]. The latter creates strong artifacts due to resize operations that directly synthesize output pixels. Our model eliminates these artifacts while using $29\%$ fewer parameters.
@@
In Fig. A15 (bottom) we remove reflections from Ansel Adams’ “Residents of Hornitos.” Ceiling light sources are visible at the top of the photo, and on the left there is a general loss of contrast. The recovered transmission image has accurately uniform contrast, and the reflection reveals the hidden image of the photographer holding their cell phone. Our model is able to recover hidden reflections like this in part due to the bit depth of RAW images. The reflection is also blurry and differs in color.
@@
Lens flares. Our model can also remove lens flares, which are reflections from the optical elements within the camera lens. An example is shown in Fig. A16. Most of the lens flare is removed, excepting one saturated region. The flare itself is also recovered in the reflection image even though the simulated reflections $r$ in the training dataset do not include lens flares. Since our method has some ability to generalize to flare removal, and it is difficult to create training data for flare removal, it could be helpful to pre-train flare removal methods first to remove photometrically accurate simulated reflections.

### v3

@@ D.1 Mixture search
Well mixed $m$ are identified by computing the SSIM between $m$ and $t$ as a block-wise image, and checking if the mean of this SSIM image is within a useful range: if the SSIM is too high, the reflection is imperceptible; if it is too low, the mixture is not visually interpretable, even by a human. We compute this single channel SSIM image as a weighted average of the corresponding per-channel SSIM images. The weights are the average value in each color channel, which better accounts for strongly colored images. Lastly, the standard deviation of this single channel SSIM image is checked to remove reflections that are imperceptible, but nonetheless produce a low mean SSIM by spreading their power broadly (they have low spatial variance). The effect of this search can be seen in Fig. S2: simulated reflections can be faint or nearly invisible (examples $19$, $22$, $23$, and $25$); or so strong that the transmission is nearly invisible or slightly difficult interpret (examples $4$, $12$, $16$, and $21$).
@@ D.2 Source images
These pairings of source images interact with the mixture search process (Sec. D.1) to introduce photometric and semantic priors, as seen in Fig. S2. Outdoor transmission scenes typically have reflections of indoor light sources or windows (examples $1$, $2$, $5$, $9$, $11$, $14$, $19$, $24$) unless it is dusk (example $3$). Indoor transmission scenes typically have strong reflections of skies or full outdoor scenes due to the brightness of natural light (examples $4$, $7$, $8$, $10$, $12$, $15$, $16$, $17$, $18$, $21$) unless it is dusk (example $13$). Lastly, indoor-indoor pairings are often complex because $t$ and $r$ are typically similar in brightness (examples $6$, $20$, $26$).
Figure S4: The base model. Mixture and context images are projected into a high dimensional space using a shared backbone [43] (labeled P-Net; weights are shared), and a feature fusion network [44] (labeled F-Net; weights are shared). The context features are used to predict affine transforms for each feature channel at each resolution. Channel-wise modulation is used because contextual photos do not always include content that can be matched. Modulation can help identify the reflection in the feature space. We use two conv-mod-deconv operations of [23] within the modulated merge blocks. The FPN combine op is a fast normalized fusion module from the BiFPN architecture [44].
@@ E.1 Base model
The base model is designed to leverage local and global features, Fig. S4, and produce $256^{2}$ pixel outputs in about $1$ second on a mobile device to meet req. 4 (see Sec. 1).
@@
The contextual image, $c$, is passed through the same F-Net and P-Net using the same weights as $m$. The features of $c$ and $m$ at the lowest resolution are input to an affine prediction module, Fig. S4 (lower right). This module first vectorizes its two $64\times 4\times 4$ inputs, and passes them through a fully connected layer to transform them into two $64$-$\mathrm{D}$ vectors. These vectors embody $64$ pairs of channels, which we concatenate and input to an MLP (labeled paired-channel MLP) that predicts affine transforms that modify the features of the FPN during the final low-to-high pass.
The paired-channel MLP is a series of grouped convolutions that implement $64$ independent MLPs followed by a fully connected layer. These $64$ MLPs each have $2$ inputs, $2$ hidden, and $1$ output dimension, with leaky ReLUs after each layer. The inputs to these MLPs are corresponding pairs of channels from $m$ and $c$. The outputs compose a single $64$-$\mathrm{D}$ vector that is input to a fully connected layer to predict $64\times K\times 2$ affine transforms, two for each of the $64\times K$ channels and levels of the FPN. [^7] Note that we include two levels at $256^{2}$ pixels, in correspondence with the resolutions of the features that we extract from the backbone. Conceptually, this paired-channel MLP has the capacity to compare $c$ and $m$ to identify channels that match the reflection scene, and to determine how to transform those channels to remove reflections.
The predicted affines from the paired-channel MLP are used in the final low-to-high pass of the FPN to perform a series of modulated merge operations, Fig. S4 (upper right). These merge operations use two affine transforms per feature channel, labeled $a$ and $b$ in Fig. S4. Transforms $a$ are used to perform a conv-mod-deconv operation ala StyleGan [23] on the features of $m$ from FPN level $K$. These features are subsequently combined with the features from level $K+1$ by resampling the latter features 2×2$\times$2 × and using fast normalized fusion [44] (labeled FPN Combine). These combined features are modified with a second conv-mod-deconv operation using the second group of affines, $b$.
The final modulated merge produces $64$-$\mathrm{D}$ features at $256^{2}$ pixels. These features are concatenated with $m$ and the features from the first layer, for a total of $3+32+64=99$ channels. A convolutional finishing module is then applied. This module has the capacity to further identify and finally render $t$ and $r$. To simplify comparison to prior work, our finishing module is the head in [59]. The first layer is $1\times 1$, and projects the $99$ input features to $64$-D, which is maintained in the remaining operations. Those operations are $3\times 3$ convolutions dilated by $(1,2,4,8,16,32,64,1)$; each is followed by a batch norm and leaky ReLU. A final $1\times 1$ convolution generates $6$ channels: $t$ and $r$.
@@ E.2 Upsampler
Figure S5: Upsampler at one pyramid level.
The upsampler is shown in Fig. S5 and introduced in Sec. 4.2. It transforms low resolution outputs $t$, $r$ from the base model to a flexible output resolution. We use a Gaussian pyramid and apply the same upsampler at each level. The upsampler projects the low-res, $3$-channel inputs $(m,r,t)$ into a high dimensional space $\phi$ using convolutions in an expand-and-contract pattern. We use $3\times 3$ kernels for feature expansion, and $1\times 1$ kernels for contraction. There are $3$ layers with hidden dimensions $(32,16),(64,32),(128,64)$. We use leaky ReLU between the hidden layers, and no skip connections. Batch norms are omitted to facilitate a feature matching process, described next.
The components $t$ and $r$ are separated by identifying low resolution features $\phi_{t}$ and $\phi_{r}$ in the low resolution mixture $\phi_{m}$. We predict $2$, per-pixel, per-channel low resolution masks using a mask prediction module, Fig. S5 (bottom), which uses a paired-channel MLP (defined in Sec. E.1) to predict its affine transforms (see also Sec. 4.2). The joint mask predictor also uses a paired-channel MLP, but it directly outputs the final masks rather than affine transforms. The input, hidden, and output dimensions of both paired-channel MLPs are $(2,2,2,1,2)$. The final layer is fully connected. As noted in Sec. E.1, these MLPs can be implemented efficiently as a series of grouped $1\times 1$ convolutions.
@@ F.1 Evaluation methods
Ground truth capture. Ground truth capture was used for Fig. 5, Fig. 8, and Fig. S6. Mixture images $m$ were captured with ground truth $r$ by placing a black material behind the glass and taking a second photo. Images $t$ were computed $t=m-r$ in linear sRGB after normalizing the exposures and using the white point of $m$ (see ACR Step 6, Sec. 3). Nonetheless, we found it necessary to capture $m$ and $r$ with fixed exposures and white points to avoid imprecisions in the values that are stored in RAW metadata.
@@
Ground truth photographs. Ground truth images were captured in three common scenarios: 1) looking out of a home window, Fig. 8; 2) photographing artwork, Fig. S6 (left); and 3) looking into a display case, Fig. S6 (right). In scenario one, the illuminant in the reflection scene is approximately $3$,$000\mathrm{K}$, and the white point of the mixture image is $7$,$300\mathrm{K}$. In scenario two, these values are $6$,$000\mathrm{K}$ and $6$,$850\mathrm{K}$, and in scenario three they are $6$,$000\mathrm{K}$ and $3$,$627\mathrm{K}$.
Figure S6: Comparisons with ground truth at $256\times 384$ and $2048\times 3072$. The re-trained, low-resolution methods Zhang [59] and DSRNet [20] are upsampled using our upsampler, and both low- and high-resolution results are shown. The SSIM of the predicted $t$ and $r$ is reported at low resolution (labeled t, r) and high (labeled T, R). Errors in ground truth are outlined and omitted from the SSIM.
@@ F.2 Base model comparisons
Reflections with ground truth. In Fig. S6 we show two images with ground truth reflections: photographing artwork, and looking into a display case. These complement Fig. 8, in which a photo was taken when looking out of a home window. The home window view includes an interior reflection that is strongly yellow due to the indoor illuminant color. In contrast, the artwork in Fig. S6 is illuminated by the same light source as the reflection scene, which produces a correctly white balanced reflection that consequently has more diverse colors. In contrast, the display case in Fig. S6 (right) reflects an outdoor scene with a different white balance, which produces a strong, blue reflection. This outdoor reflection is visible over broad regions because the illuminant is powerful enough to reflect off of the diffuse ground and sidewalk surfaces with sufficient intensity to be visible over the contents of the display case. These exterior reflections are also qualitatively different from those in photo of the artwork, where the reflection is sparse. The SSIM of the artwork is therefore high on average ($0.994$), but the reflections are locally strong, whereas the SSIM of the display case is low ($0.833$) because the reflection affects broad regions.
@@
Our base models improve the SSIM of $t$ and $r$ in all of these ground truth examples (labeled in lower case t, r), and this extends to the upsampled results (labeled in uppercase T, R) whereas prior works do not perform as well (Fig. S6 and Fig. 8). Our contextual model produces a more correct transmission and reflection image on the artwork. On the display case, the contextual model improves the reflection, whereas the cars are not fully removed. We believe this variation results from saturated regions in the sky of the contextual photo, where we use the as-shot illuminant color in the EXIF to recover the color of the saturated pixels. Lastly, the method of Zhang [59] associates blue colored content with the reflection in both images, but this is incorrect for the painting; the transmission image is therefore distorted.
@@
Figure S7: Results for base models at $256^{2}$ pixels. Reflections marked “+X” are brightened by X stops compared to the transmission.
@@
Figure S8: Results for base models at $256^{2}$ pixels. Reflections marked “+X” are brightened by X stops compared to the transmission.
Figure S9: Reflection recovery. Our model separates reflections that can be difficult to spot with the naked eye. See Fig. S11.
@@
We found that our model removes blue colored reflections consistently well, and we believe that this results from their commonness (outdoor illuminants are powerful and therefore frequently create reflections that appear blue when mixed with interior scenes). The yellow color of interior reflections on outdoor scenes also seems to help, as our model can separate textured objects like the painting on the wall in Fig. 8 from the tree texture. We also tried illuminating the indoor scene in Fig. 8 with a special studio light that is much more powerful than a typical interior light source. This created a warm indoor reflection that was analogous to the display case in Fig. S6 (right), where the reflection covers large parts of the transmission scene. Our model results are less consistent in this artificial situation. We believe this is because it is rare for interiors to be flooded with such strong lights, and so their reflections are uncommon in our training data. We find that interior reflections tend to appear in small regions because most artificial light sources are weak—only objects near the light will be bright enough to reflect over outdoor scenes. At nighttime, however, consumer illuminants easily reflect over dark cityscapes. We found that our model results are less consistent on such images. This can be improved by augmenting the dataset with source images $t=T(i)$ that were taken at nighttime.
Figure S10: Removing lens flares. Although our training data do not include images of lens flares, our model can sometimes remove them.
Figure S11: Upsampler performance comparison. We upsample a ground truth reflection image from $256\times 384$ to $2048\times 3072$ using our method and V-DESIRR [35]. The latter creates strong artifacts due to resize operations that directly synthesize output pixels. Our model eliminates these artifacts while using $29\%$ fewer parameters.
@@
Reflections in the wild. In Fig. S7, S8, and Fig. 8 we show results on reflections that were captured in-the-wild, where it was not possible to capture ground truth: (left) shopping in the morning, (middle) looking into a building at midday in a city, and (right) photographing artwork in an outdoor mall (Fig. S7); while traveling and photographing artwork from a city street (S8). We compare to Zhang [59] and DSRNet [20] using their published $8$-bit models (bottom two rows). The $8$-bit models do not consistently remove in-the-wild reflections, with the exception that the method of Zhang et al. seems to have learned to remove blue colored content (left); see also Fig. S9 (top). We retrained these prior works on our photometrically accurate training data (marked “retrained”), and they improve significantly. This suggests that the muted response of the $8$-bit models on in-the-wild reflections results from differences between their training data and real world reflections, and furthermore that the training process of prior works is insufficient: pre-training on photometrically inaccurate images, and fine-tuning on small datasets of ground truth reflections does not produce models that generalize as well. Looking closely, however, note that the retrained models do still introduce blur and colored artifacts.
@@
Our methods perform well across these diverse in-the-wild use cases. In Fig. S7, both our contextual and non-contextual models separate the strongly blue colored reflection, the complex cityscape reflection, and the artwork reflection (in the entrance to an indoor shopping mall). The contextual model improves each of these cases, excepting the transmission image for the shopping photo. We believe this also results from the saturated regions in the upper right of the contextual photo, where we have used the as-shot illuminant color in the EXIF to recover the pixels.
@@
Subtle reflections. In Fig. S9 we show results on images with subtle reflections, which are common in art museums where the glass and lighting are optimized to reduce reflections. Our models are able to remove these reflections, and recover $r$ even when it is invisible to the naked eye. In the painting (Rembrandt’s “St. Matthew and the Angel”), the recovered reflection has a strong blue color due to the special glass that is used in museums, and it correctly depicts the photographer and gallery. The colors in the recovered painting are also more correct. The specular reflection from the surface of the painting (top) is not removed, and we believe that this is the desired result. We have found that our models will sometimes remove sharp specular regions that are similar in color to the reflection from the glass, and this is visible in the upper quarter of the reflection image, where some specular texture from the surface of the painting is visible, and has been removed from the transmission.
@@
In Fig. S9 (bottom) we remove reflections from Ansel Adams’ “Residents of Hornitos.” Ceiling light sources are visible at the top of the photo, and on the left there is a general loss of contrast. The recovered transmission image has accurately uniform contrast, and the reflection reveals the hidden image of the photographer holding their cell phone. Our model is able to recover hidden reflections like this in part due to the bit depth of RAW images. The reflection is also blurry and differs in color.
Our ability to uncover subtle reflections is in part due to the extended bit depth of RAW images. The images in Fig. 1, Fig. 5-8, Fig. S6-S15 are $\geq$15-bit. ACR Step 3 is uint16. Together, the source datasets for the simulation (MIT5K [8] and RAISE [10]), are 12- and 14-bit (43%, 57%).
@@
Lens flares. Our model can also remove lens flares, which are reflections from the optical elements within the camera lens. An example is shown in Fig. S10. Most of the lens flare is removed, excepting one saturated region. The flare itself is also recovered in the reflection image even though the simulated reflections $r$ in the training dataset do not include lens flares. Since our method has some ability to generalize to flare removal, and it is difficult to create training data for flare removal, it could be helpful to pre-train flare removal methods first to remove photometrically accurate simulated reflections.
Failures. Failures typically occur when an image is too difficult for a person to visually interpret, as shown in Fig. S12. Light sources might also be missed, or a halo might be left behind, but in-painting methods can fill these holes. Reflections are usually removed best when glass covers the field-of-view, blocking the photographer from their subject, as this matches how the simulation is designed. As a result, local reflections on objects are typically not modified, which is desirable (e.g. shiny cars, distant windows).
Results on 8-bit photos. In this work, all results are presented on RAW photos because our models are trained for RAW. Nonetheless, our RAW-trained models have some ability to remove reflections in 8-bit photos if gamma is pseudo-inverted, as shown in Fig. S13, but removal may be incomplete. This difference between RAW and 8-bit photos appears because the latter are produced by diverse camera pipelines and artistic adjustments. This diversity of finishing operations must be considered when training and testing models that have been trained on RAW-based simulation images. It is therefore not meaningful to test our models on datasets of 8-bit images.
Figure S12: Failure cases. Hard-to-interpret photos can cause reflection removal to be imperfect.
Figure S13: Results when running a RAW-trained model on 8-bit inputs. Results on 8-bit photos can be obtained by pseudo-inverting gamma, but they are often imperfect because such photos have undergone a variety of non-linear finishing effects that are not present in RAW images (see ACR Steps 7-8).
Figure S14: Upsampling errors. V-DESIRR [35] re-introduces low-frequency reflections. Our method avoids this error, but copies high-frequency textures that were not visible at $256^{2}$.
Figure S15: Predicting the reflection enables aesthetic editing and error correction.